OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? Answer 107π in3 102π in3 106π in3 108π in3

OpenStudy (cwrw238):

volume of water = volume of empty vase - total volume of the 6 marbles = pi*3^2 * 12 - 6* (4/3) pi *1^3 can you do this calculation?

OpenStudy (anonymous):

i beleive so hold on

OpenStudy (cwrw238):

i make it 100 pi - either i've gone wrong or the choices are incorrect!!

OpenStudy (anonymous):

aha i got 173.6 .....

OpenStudy (cwrw238):

i can't see how i've gone wrong - the vol of cylinder is area of saw * ht = pir^2 * h = pi *3^3 * 12 = 108 pi

OpenStudy (anonymous):

108 is one of my options

OpenStudy (cwrw238):

- thats the total volume - if we got 6 marbles of volume 4/3 pi volume = 9pi that leaves 100pi of water

OpenStudy (anonymous):

idk the smallest i have is 102 so

OpenStudy (cwrw238):

sorry 8pi so water vol. = 108pi - 8pi = 100pi

OpenStudy (cwrw238):

well a marble is a sphere whose volume is (4/3) pi r^3 and r = 1 (as diameter = 2) so that 6* 4/3 pi = 8 pi

OpenStudy (cwrw238):

one of the options should be 100pi

OpenStudy (anonymous):

=/

OpenStudy (cwrw238):

ok - i'm sure of that (hope i don't sound arrogant) - whats =/ mean?

OpenStudy (anonymous):

no it's fine the options could be wrong i mean they have messed up before so .. yeahh and its a straight face lol

OpenStudy (cwrw238):

ok - lol